4 Consider a system with 14 dedicated devices of the same type. All jobs currently running on this system require a maximum of five devices to complete their execution but they each run for long periods of time with just three devices and request the remaining two only at the very end of the run. Assume that the job stream is endless and that your operating system’s device allocation policy is a very conservative one: No job will be started unless all the required drives have been allocated to it for the entire duration of its run. a. What is the maximum number of jobs that can be active at once? b. What are the minimum and maximum number of devices that may be idle because of this policy? Under what circumstances would an additional job be started?



1
Expert's answer
2021-11-16T18:36:03-0500

Since a very conservative allocation policy is used jobs can only be started when the required number of devices are available.

a. maximum number of jobs that can be active be x.

5*x<=14 i.e x=2 hence maximum number of jobs that can be active =2.

b. Minimum number of idle devices are present when both the jobs are using 5 devices = 14 - 2*5 = 4.

Maximum number of idle devices are present when jobs are using 3 devices = 14 - 2*3 =8.